OpenStudy (maths99):

Im having problems with uncertainty. Please help. Complete the table for 1/m^2 and put its uncertainty.

OpenStudy (maths99):



ganeshie8 (ganeshie8):

May I suggest you go through this first http://web.uvic.ca/~jalexndr/192UncertRules.pdf

OpenStudy (maths99):

Okay ive already read that, but i still need help for this uncertainty.

OpenStudy (irishboy123):

you seem to know some calculus so this is a quick way to do it  alternatively, you will calculate \(\dfrac{1}{(m-0.01)^2}\) and \(\dfrac{1}{(m+0.01)^2}\) was well as \(\dfrac{1}{m^2}\) and create a new range of values for that particular \(\dfrac{1}{m^2}\). plus there's some round ing to do.